# for the Baryon Scattering (BaSc) Collaboration

### v1

@@ III.2 Off diagonal correlators with HX operators
The discrepancy boils down to a difference in energy extractions. Effective masses from Ref. Beane et al. (2017) for the deuteron (lower left plot) and dineutron (lower right plot) are shown in Fig. 14 for a $48^{3}$ spatial lattice. A single off-diagonal correlator involving a local hexaquark operator at the source and a nucleon-nucleon operator of zero momentum at the sink was used to determine these effective masses, which are shown as blue circles with errors. The horizontal green lines indicate the approximate energy extractions from Ref. Beane et al. (2013a) with the vertical thickness indicating the statistical uncertainties. The horizontal dashed lines indicate the energies of two non-interacting nucleons at rest in the deuteron and dineutron cases. The occurrences of the horizontal green lines well below the horizontal dashed lines, along with similar differences in other energy extractions, essentially lead to the bound states observed, for example, in the upper plot. The horizontal violet boxes indicate approximately where the energy extractions from Ref. Hörz et al. (2021) would be for this $48^{3}$ lattice. The lowest-lying energies in Ref. Hörz et al. (2021) occur very slightly below the non-interacting energies. Hence, the discrepancy essentially arises from the differences between the green and violet lines, along with similar differences from other energy determinations.
Figure 15: (Upper left) Same as the bottom left plot of Fig. 14, but also includes the fit from Eqs. (29), (30), (31) as a violet curve. For temporal range $t/a=10-20$, the violet curve shows a remarkable but false plateau. (Upper right) Same as the upper left plot in this figure, but continued to much larger temporal separations, showing the very slow approach to the asymptotic limit. (Lower left) Same as the bottom right plot of Fig. 14, but also includes the fit from Eqs. (29), (30), (32) as a violet curve. (Lower right) Same as the lower left plot in this figure, but continued to much larger temporal separations.
@@ IV HAL QCD Potential
Numerical determination of the potential, $U(\mathbf{r},\mathbf{r^{\prime}})$ in general, is challenging. In practice, in both the literature and in this work, the leading local approximation of this non-local potential is used:
@@
where additional terms include gradient terms as well as spin-orbit mixing. While such a gradient expansion may suggest some sort of low-energy approximation, there is no known relevant energy scale to properly cut off such an expansion, which is, in general, uncontrolled. Possibilities for testing the relative sizes of such corrections have been proposed by the HAL QCD collaboration, and involve the use of two-nucleon correlation functions having different source operators Iritani et al. (2019b). In such cases it is imperative to fully control the large-time behavior of the calculated potential for each source operator. We will explore the time-dependence in detail in this work.
@@
Once the (local) potential has been determined numerically, scattering phase shifts are extracted through solution to the Schrödinger equation, typically first by fitting the numerical potential data to analytic functions. The HAL QCD collaboration has largely utilized a Gaussian basis to describe the potential Ishii et al. (2012), but they have also explored the possibility of adding a one- and two-pion exchange term Iritani et al. (2019b). In this work, we thoroughly investigate the effects of different analytic forms used to fit the potential, including incorporating regulated one- and multiple-pion exchange potentials, and exploring the use of Harmonic Oscillator basis functions as well as Gaussians.
@@
In addition to the aforementioned tests of systematics of the HAL potential, we furthermore explore the effects of the finite-volume on the extracted potential, and the effects of discretization of the NBS equation. The results of these studies are incorporated into an overall systematic error on the extracted scattering phase shifts through a bootstrap analysis. Details of these systematics are discussed below.
@@ IV.1.1 Excited state contamination
where ${\Delta}^{\rm NN,I}_{1}$ is the mass gap from the ground state NN energy to the first inelastic NN state and ${\Delta}^{\rm N}_{1}$ is the mass gap to the first single nucleon inelastic mass gap. See Appendix B.1 for a derivation of this formula.
@@
In Fig. 18, we show the potential in the deuteron channel for various Euclidean times, as well as the extrapolation of the potential to infinite time. The same plot for the di-neutron channel is presented in Appendix B.3. Performing such extrapolations is clearly merited for controlling systematics in precision calculations utilizing the potential method, as choosing a single fixed value of the time, $t$, may be biased due to correlated fluctuations. Furthermore, due to the $\mathbf{r}$-dependence of the overlap onto excited states discussed previously, it may not be possible to choose a single value of $t$ such that inelastic excited states are well-controlled for all values of $r$. In Fig. 19, we show the resulting phase shift as a function of scattering momentum squared utilizing a given fixed value of $t$. Different forms of the short-distance portion of the fit function, denoted “harmosc” (“gaussian”) for the Harmonic Oscillator (Gaussian) basis. We find that for early and intermediate times, particularly in the deuteron channel, the different choices for the form short-distance potential give results that disagree significantly from each other, even for very low energies (large $\mathbf{r}$), signaling combined discretization/excited state contaminations. For later times and for the extrapolated potential, we find no significant differences between the different fit forms and include both in our quoted systematic error band.
Figure 19: Scattering phase shift in the deuteron channel for various Euclidean time, $t$, as a function of momentum transfer in the center of mass frame, $q_{\rm CM}$, in units of the pion mass, $m_{\pi}$. Phase shifts are determined using two possible fit functions, having a Harmonic Oscillator (“harmosc”) or Gaussian (“gaussian”) basis for the short-range portion of the potential.
@@ IV.1.2 Discretization effects
and $\nabla_{a}$ is any discretized version of the Laplacian. We have used the symmetric definition for the temporal derivative, such that the leading lattice spacing dependence from this term comes in at $\mathcal{O}(a^{2})$.
@@
Finally, we test systematics coming from the short-range portion of the potential through various tests of the fit model. The HAL QCD collaboration has utilized a Gaussian basis to describe the potential as well as exploring the possibility of adding a one-pion exchange term for the long-range tail and rho-exchange for intermediate range contributions Aoki et al. (2010). In this work, we thoroughly investigate the effects of different analytic forms used to fit the potential, including incorporating regulated one- and multiple-pion exchange potentials (see Sec. IV.1.3), and exploring the use of Harmonic Oscillator basis functions versus Gaussians. With differences in the results from different bases for the short-range potential, and variation in $r_{\rm min}$ included in the fit, we can probe discretization effects coming from short-range contributions. Fits utilizing $r_{\textrm{min}}\lesssim 0.5$ fm and both Harmonic Oscillator and Gaussian fits are included in our final quoted systematic.
Another discretization systematic arises from the modified dispersion relation, Eq. (17) for which we found $\xi=0.979(5)$, Eq. (19). A proposal to account for this effect is to modify Eq. (36) by replacing $\nabla^{2}\rightarrow\xi\nabla^{2}$, see App. B.2. As with the other discretization effects discussed in this section, we find that this does not have a statistically significant effect on the potential or subsequent value of $q\cot{\delta}$.
@@ IV.1.3 Finite volume effects
We test sensitivity to finite volume effects in multiple ways. First, we test different forms for the long-range terms in the fit function to the potential. For this, we use a regulated Yukawa potential
$V^{(p)}(r)=\frac{e^{-mr}}{mr}\left(1-e^{-cr^{2}}\right)^{p}\ ,$ (42)
where $m,c,p$ are free parameters. Parameter $m$ models the mass of the exchange particle. Parameters $c$ and $p$ control a regulator that removes the non-physical short-range divergence that comes from treating nucleons as point particles. The regulator is borrowed from the Argonne $v_{18}$ one pion exchange potential implementation Wiringa et al. (1995), equations 17 through 19, with $c$ (related to the nucleon size) controlling the range of regulator and $p$ its strength. We performed fits with $p=1$ and $p=2$, with results favoring the more rapid cutoff with $p=2$. We find no sensitivity to the number of terms (representing multi-meson fits) included in the fit function, and also find that the fit is fairly insensitive to any of these fit parameters. We furthermore test sensitivity of the fit to different cutoffs at large $r>r_{\mathrm{max}}$. The fits are again nearly completely insensitive to these changes, as shown in Fig. 22. We include fits using $r_{\mathrm{max}}\gtrsim 1.3$ fm as part of our quoted systematic band.
@@ IV.2 Final results
In Fig. 23, we show the potential extrapolated to $t\to\infty$ as a function of $r$ (black data). Fits to this data that are included in our final error band for the phase shift are shown as blue (purple) bands, representing fits utilizing both single- and double-meson exchange, Eq. (42), and a sum of 2-3 Gaussians (Harmonic Oscillator functions) to represent the short-range portion of the potential. While the fits using Gaussians and Harmonic Oscillator functions show completely different behavior for $r$ smaller than the included fit range, we find no significant difference in the resulting phase shifts within the energy range of interest (Fig. 23 (right)). We therefore conclude that the physics of interest is insensitive to the behavior of the potential below this range. We also find no significant differences in our fitted potential having greater than 1 long-range term and greater than 2 short-range terms. The chosen fits furthermore include two different long-range cutoffs, $r_{\mathrm{max}}\sim 1.3,1.45$ fm, and two different short-range cutoffs, $r_{\textrm{min}}\sim 0.4,0.5$ fm, for which we saw relative stability in the phase shift results (see previous Sections).
@@
Figure 23: Potential, $V_{C}(\mathbf{r})$ (left), in the deuteron channel as a function of $r$, in units of fm (black data). Colored bands represent all fits included in the final analysis, where blue (purple) bands utilized a sum of Gaussians (Harmonic Oscillator functions) as a basis for the short-range contributions. Vertical dashed lines represent the two possible minimum values of $r$ included in the fits. The right figure shows the resulting $q\cot{\delta}$ arising from these two different choices of the short-range potential over the full range of fits explored. One can see that the choice of short-range potential and the choice of $r_{\rm min}$ do not have any significant impact on the determination of $q\cot{\delta}$ in this low-energy regime.
@@ V Discussion and conclusion
We presented a high-statistics calculation of the NN scattering amplitudes in the SU(3) flavor limit with a heavy pion mass of $m_{\pi}\approx 714$ MeV. The calculation was designed to target a mass region where a discrepancy in the literature on the nature of the NN systems, whether or not they form bound states, has persisted for more than a decade. In order to address this controversy, for the first time, we computed the NN amplitudes with several methods in the literature on the same set of underlying gauge ensembles, allowing us to isolate systematic uncertainties arising in the methods.
@@
In Fig. 25, we plot the resulting $q\cot{\delta}$ results from both the single channel finite-volume QC2 analysis as well as from the HAL QCD potential, which show qualitative agreement between the methods. The source of small quantitative discrepancy between these methods is not currently resolved. As discussed in the present work, the discretization errors between the methods are not the same so a continuum extrapolation is required for such a comparison. If discretization effects are the source of discrepancy, those in Eq. (IV.1.2) seem the most likely as they are only exponentially suppressed by energy gaps to the elastic NN scattering modes, and thus both difficult to resolve and more relevant for low-energy. Another obvious candidate to check is the derivative expansion of Eq. (37), which we have not explored in this work. Despite this small discrepancy, both methods exclude bound states.
@@
In Sec. III.2, we discussed the toy-model presented by NPLQCD in Ref. Amarasinghe et al. (2023) and in Chapter 15 of Ref. Tews et al. (2022) that argues the off-diagonal correlators may be able to capture a deeply-bound state that the other operators are not sensitive to. Recognizing the unrealistic and contrived nature of their model, we made small, simple and physically reasonable changes to the model and showed that the off-diagonal hexa-quark correlator produced no advantage whatsoever.
We conclude that di-nucleons do not form bound states at heavy pion mass. In the case of the deuteron, the QC2 analysis rules out a bound state on this ensemble by $>5\sigma$. In the case of the di-neutron, a bound state is excluded at the $\lesssim 3{\sigma}$ level. Previous identification of bound and deeply-bound di-nucleons Beane et al. (2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Wagman et al. (2017) most likely arose from a misidentification of a plateau in the effective energy arising from opposite sign contributions allowed in off-diagonal correlation functions, as first suggested by HAL QCD Iritani et al. (2016). This observation also invalidates all two- and higher-body LQCD matrix element calculations that utilized these off-diagonal correlation functions as those results relied upon the assumption of deeply dound di-nucleons to relate the finite and infinite volume matrix elements, which are in general non-trivial in magnitude, see for example Briceno et al. (2015), thus leaving the previous results with an unquantified systematic.

### v2

@@ III.2 Off diagonal correlators with HX operators
The discrepancy boils down to a difference in energy extractions. Effective masses from Ref. Beane et al. (2017) for the deuteron (lower left plot) and dineutron (lower right plot) are shown in Fig. 15 for a $48^{3}$ spatial lattice. A single off-diagonal correlator involving a local hexaquark operator at the source and a nucleon-nucleon operator of zero momentum at the sink was used to determine these effective masses, which are shown as blue circles with errors. The horizontal green lines indicate the approximate energy extractions from Ref. Beane et al. (2013a) with the vertical thickness indicating the statistical uncertainties. The horizontal dashed lines indicate the energies of two non-interacting nucleons at rest in the deuteron and dineutron cases. The occurrences of the horizontal green lines well below the horizontal dashed lines, along with similar differences in other energy extractions, essentially lead to the bound states observed, for example, in the upper plot. The horizontal violet boxes indicate approximately where the energy extractions from Ref. Hörz et al. (2021) would be for this $48^{3}$ lattice. The lowest-lying energies in Ref. Hörz et al. (2021) occur slightly below the non-interacting energies. Hence, the discrepancy essentially arises from the differences between the green and violet lines, along with similar differences from other energy determinations.
Figure 16: (Upper left) Same as the bottom left plot of Fig. 15, but also includes the fit from Eqs. (31), (32), (33) as a violet curve. For temporal range t/a=1020t/a=10-20, the violet curve shows a remarkable but false plateau. (Upper right) Same as the upper left plot in this figure, but continued to much larger temporal separations, showing the slow approach to the asymptotic limit. (Lower left) Same as the bottom right plot of Fig. 15, but also includes the fit from Eqs. (31), (32), (34) as a violet curve. (Lower right) Same as the lower left plot in this figure, but continued to much larger temporal separations.
@@ IV HAL QCD Potential
Numerical determination of the potential, $U(\mathbf{r},\mathbf{r^{\prime}})$ in general, is challenging. In practice, in both the literature and in this work, the leading local approximation of this non-local potential is most commonly used:
@@
where additional terms include gradient terms as well as spin-orbit mixing. HAL QCD has investigated both of these extra contributions Murano et al. (2014); Iritani et al. (2019b). In this work we neglect higher order terms in the gradient expansion. Further, we restrict our attention to the $T_{1g}$ ($A_{1g}$) irreps for the deuteron (di-neutron) which are dominated by the central potential $V_{C}(r)$ at low energy. As discussed in App. B.4, we verified that remaining $\ell\geq 4$ contributions are negligible by performing an $\ell=0$ projection of the NN correlator data and comparing the resulting potential and phase shifts. This is similar to the use of the Misner method Misner (2004) by the HAL QCD collaboration Miyamoto et al. (2020).
@@
While such a gradient expansion may suggest some sort of low-energy approximation, there is no known relevant energy scale to properly cut off such an expansion, which is, in general, uncontrolled. Possibilities for testing the relative sizes of such corrections have been proposed by the HAL QCD collaboration, and involve the use of two-nucleon correlation functions having different source operators Iritani et al. (2019b). In such cases it is imperative to fully control the large-time behavior of the calculated potential for each source operator. We will explore the time-dependence in detail in this work.
@@
In this work, we investigate the effects of different analytic forms used to fit the potential, including incorporating regulated one- and multiple-pion exchange potentials, and exploring the use of harmonic oscillator (H.O.) basis functions as well as Gaussians for the short range contribution $V_{s}(r)$. The HAL QCD collaboration has primarily used a Gaussian basis to describe the potential Ishii et al. (2012), but they have also explored the possibility of adding a one- and two-pion exchange terms Iritani et al. (2019b). The use of harmonic oscillator basis functions is new in this work. The various potentials we fit to the results include
$V(r)=V_{s}(r)+V_{\pi}(r)\;\left[\;+\;V_{2\pi}(r)\;\right]\,$, (40)
where the $V_{2\pi}(r)]$ indicates it is optionally included.
For the one- and two-pion exchange potentials, we utilize a regulated Yukawa potential of the form
$V_{[2]\pi}^{(p)}(r)=-g^{2}\frac{e^{-mr}}{r}\left(1-e^{-cr^{2}}\right)^{p}\ ,$ (41)
where $m,c,p$ are free parameters. The coupling is modeled as $g^{2}=e^{a}$ and we have absorbed the $m^{-1}$ that usually accompanies the $r^{-1}$. Parameter $m$ models the mass of the exchange particle. Parameters $c$ and $p$ control a regulator that removes the non-physical short-range divergence that comes from treating nucleons as point particles. The regulator is borrowed from the Argonne $v_{18}$ one pion exchange potential implementation Wiringa et al. (1995), equations 17 through 19, with $c$ (related to the nucleon size) controlling the range of regulator and $p$ its strength. We performed fits with $p=1$ and $p=2$, with results favoring the more rapid cutoff with $p=2$. We find no sensitivity to the number of terms (representing multi-meson fits) included in the fit function, and also find that the fit is fairly insensitive to any of these fit parameters.
For the short-range terms, we explore three different functional forms
$V_{s}(r)=\left\{\begin{matrix}\sum_{i}v_{i}e^{-(r/b_{i})^{2}}&\textrm{``Gaussian"}\\ \sum_{i}v_{i}e^{-1/2(r/b)^{2}}L^{1/2}_{i-1}\left(\frac{r^{2}}{b^{2}}\right)&\textrm{``harmosc"}\\ \sum_{i}v_{i}e^{-(r/b_{i})^{2}}H_{i}(r/b_{i})&\textrm{``1Dosc"}\end{matrix}\right$.\,, (42)
where $v_{i}$ are strengths for the various terms. For all three functions, the $b_{i}$ are length scales associated with different order terms. In the case of the 3D harmonic oscillator, the $b$ parameter is the same for all terms to preserve the orthogonality of the basis functions. $L_{i}$ and $H_{i}$ are Laguerre and Hermite polynomials respectively. The Gaussian basis is what is often used by HAL QCD. The 3D Harmonic oscillator choice is a complete orthogonal basis commonly used on calculations of nuclei with nucleon degrees of freedom. The 1D oscillator functions are used to demonstrate that the functional form used to describe the potential has little impact on the resulting phase shift, provided the basis is able to describe the numerical potential data in the relevant region of $r$ that effects the phase shift at low energies.
Given a parameterization of the potential, with parameters constrained from the results, we investigate extracting the $S$-matrix in two different ways, the first of which is the standard method of numerically solving the Schrödinger equation to determine the asymptotic wavefunction. The second uses the variable phase method Calogero (1968), see Chapter 19, equation 19. A simple first-order differential equation for $S(R)$ involving $V(r)\Theta(R-r)$ is given, where the Heaviside function forces the potential to 0 outside of range $R$. The initial condition is $S(0)=I$, and $S(\infty)=S$, the desired $S$-matrix. One advantage with this method for future work is that coupled channels are directly supported. Also see Das (1981) for an extension to non-local potentials.
@@
In addition to the aforementioned tests of systematics of the HAL QCD potential, we also explore the effects of the finite-volume on the extracted potential, and the effects of discretization of the NBS equation. The results of these studies are incorporated into an overall systematic error on the extracted scattering phase shifts through a bootstrap analysis. Details of these systematics are discussed below.
@@ IV.1.1 Excited state contamination
where ${\Delta}^{\rm NN,I}_{1}$ is the mass gap from the ground state NN energy to the first inelastic NN state and ${\Delta}^{\rm N}_{1}$ is the mass gap to the first single nucleon inelastic mass gap. See Appendix B.1 for a derivation of this formula. This motivates a multi-exponential fit of the form
$V_{C}(t,\mathbf{r})=V_{C}(\mathbf{r})+\sum_{n=1}\alpha_{n}(\mathbf{r})e^{-{\Delta}_{n}(\mathbf{r})t}\,$, (45)
where $\alpha_{n}(\mathbf{r})$ and ${\Delta}_{n}(\mathbf{r})$ are independently fit for each $\mathbf{r}$.
@@
In Fig. 19, we show the potential in the deuteron channel for various Euclidean times, as well as the extrapolation of the potential to infinite time. The same plot for the di-neutron channel is presented in Appendix B.3. Performing such extrapolations is clearly merited for controlling systematics in precision calculations utilizing the potential method, as choosing a single fixed value of the time, $t$, may be biased due to correlated fluctuations. Furthermore, due to the $\mathbf{r}$-dependence of the overlap onto excited states discussed previously, it may not be possible to choose a single value of $t$ such that inelastic excited states are well-controlled for all values of $r$. In Fig. 20, we show the resulting phase shift as a function of scattering momentum squared utilizing a given fixed value of $t$. Different forms of the short-distance portion of the fit function, denoted “harmosc” (“gaussian”) for the Harmonic Oscillator (Gaussian) basis. We find that for early and intermediate times, particularly in the deuteron channel, the different choices for the form of the short-distance potential give results that disagree significantly from each other, even for low energies (large $\mathbf{r}$), signaling combined discretization/excited state contaminations. For later times and for the extrapolated potential, we find no significant differences between the different fit forms and include both in our quoted systematic error band.
Figure 20: Scattering phase shift in the deuteron channel for various Euclidean time, $t$, as a function of momentum transfer in the center of mass frame, $q_{\rm CM}$, in units of the pion mass, $m_{\pi}$. Phase shifts are determined using two examples of our fit functions from Eq. 42, 1Dosc (solid) or gaussian (dashed), for the short-range portion of the potential.
@@ IV.1.2 Discretization effects
and $\nabla_{a}$ is any discretized version of the Laplacian. We have used the symmetric definition for the temporal derivative, such that the leading lattice spacing dependence from this term comes in beyond $\mathcal{O}(a^{2})$.
@@
Finally, we test systematics coming from the short-range portion of the potential through various tests of the fit model. The HAL QCD collaboration has utilized a Gaussian basis to describe the potential as well as exploring the possibility of adding a one-pion exchange term for the long-range tail and rho-exchange for intermediate range contributions Aoki et al. (2010). In this work, we thoroughly investigate the effects of different analytic forms used to fit the potential, including incorporating regulated one- and multiple-pion exchange potentials (see Sec. IV.1.3), and exploring the use of Harmonic Oscillator basis functions versus Gaussians. With differences in the results from different bases for the short-range potential, and variation in $r_{\rm min}$ included in the fit, we can probe discretization effects coming from short-range contributions. Fits utilizing $r_{\textrm{min}}\lesssim 0.5$ fm and all three forms for the short-distance fit function are included in our final quoted systematic.
Another discretization systematic arises from the modified dispersion relation, Eq. (18) for which we found $\xi=0.979(5)$, Eq. (20). A proposal to account for this effect is to modify Eq. (38) by replacing $\nabla^{2}\rightarrow\xi\nabla^{2}$, see App. B.2. As with the other discretization effects discussed in this section, we find that this does not have a statistically significant effect on the potential or subsequent value of $q\cot{\delta}$. Since HAL QCD noted some impact of higher $\ell\geq 4$ contributions that they removed with an $\ell=0$ projection using the Misner method Misner (2004), we also performed an $\ell=0$ projection, finding that the $\ell\geq 4$ contributions are sub-percent variations that appear as a small noise contribution and do not affect our phase shift results, see App. B.4.
@@ IV.1.3 Finite volume effects
We test sensitivity to finite volume effects in multiple ways. First, we test different forms for the long-range terms in the fit function to the potential. For this, we use a sum of regulated Yukawa potentials, varying the number of included terms Eq. (41). We furthermore test sensitivity of the fit to different cutoffs at large $r>r_{\mathrm{max}}$. The fits are again nearly completely insensitive to these changes, as shown in Fig. 23. We include fits using $r_{\mathrm{max}}\gtrsim 1.3$ fm as part of our quoted systematic band.
@@ IV.2 Final results
In Fig. 24, we show the potential extrapolated to $t\to\infty$ as a function of $r$ (black data). Fits to this data that are included in our final error band for the phase shift are shown as blue (purple) bands, representing fits utilizing both single- and double-meson exchange, Eq. (41), and a sum of 2-3 Gaussians (Harmonic Oscillator functions) to represent the short-range portion of the potential. While the fits using Gaussians and Harmonic Oscillator functions show completely different behavior for $r$ smaller than the included fit range, we find no significant difference in the resulting phase shifts within the energy range of interest (Fig. 24 (right)). This also means the phase shift in this energy range is not sensitive to the divergent nature of the short range part of the potential Aoki et al. (2013). We therefore conclude that the physics of interest is insensitive to the behavior of the potential below this range. We also find no significant differences in our fitted potential having greater than 1 long-range term and greater than 2 short-range terms. The chosen fits furthermore include two different long-range cutoffs, $r_{\mathrm{max}}\sim 1.3,1.45$ fm, and two different short-range cutoffs, $r_{\textrm{min}}\sim 0.4,0.5$ fm, for which we saw relative stability in the phase shift results (see previous Sections).
@@
Figure 24: Potential, $V_{C}(\mathbf{r})$ (left), in the deuteron channel as a function of $r$, in units of fm (black data). Colored bands represent all fits included in the final analysis, where the different colored bands utilize different functions as a basis for the short-range contributions (Eq. 42). Vertical dashed lines represent the two possible minimum values of $r$ included in the fits. The right figure shows the resulting $q\cot{\delta}$ arising from these two different choices of the short-range potential over the full range of fits explored. One can see that the choice of short-range potential and the choice of $r_{\rm min}$ do not have any significant impact on the determination of $q\cot{\delta}$ in this low-energy regime.
@@ V Discussion and conclusion
We presented a high-statistics calculation of the NN scattering amplitudes in the SU(3) flavor limit with a heavy pion mass of $m_{\pi}\approx 714$ MeV. The calculation was designed to target a mass region where a discrepancy in the literature on the nature of the NN systems, whether or not they form bound states, has persisted for more than a decade, see for example the Lattice 2012 and 2013 reviews Doi (2012); Walker-Loud (2014). In order to address this controversy, for the first time, we computed the NN amplitudes with several methods in the literature on the same set of underlying gauge ensembles, allowing us to isolate systematic uncertainties arising in the methods.
@@
In Fig. 26, we plot the resulting $q\cot{\delta}$ results from both the single channel finite-volume QC2 analysis as well as from the HAL QCD potential, which show qualitative agreement between the methods. The source of small quantitative discrepancy between these methods is not currently resolved. As discussed in the present work, the discretization errors between the methods are not the same so a continuum extrapolation is required for such a comparison. If discretization effects are the source of discrepancy, those in Eq. (IV.1.2) seem the most likely as they are only exponentially suppressed by energy gaps to the elastic NN scattering modes, and thus both difficult to resolve and more relevant for low-energy. Another obvious candidate to check is the derivative expansion of Eq. (39), which we have not explored in this work. Finally, it would be interesting to use the HAL QCD method to construct optimized operators Lyu et al. (2025) and then use the correlation functions built from them to see if improvements in the extraction of the energy and subsequent QC2 can be obtained. Despite this small discrepancy, both methods exclude bound states.
@@
In Sec. III.2, we discussed the toy-model presented by in Ref. Amarasinghe et al. (2023) and in Chapter 15 of Ref. Tews et al. (2022) that argues the off-diagonal correlators may be able to capture a deeply-bound state that the other operators are not sensitive to. Recognizing the unrealistic and contrived nature of their model, we made small, simple and physically reasonable changes to the model and showed that the off-diagonal hexa-quark correlator produced no advantage whatsoever.
The present work utilized a single lattice spacing and also a different clover-action from that in Refs. Beane et al. (2013a); Berkowitz et al. (2017); Wagman et al. (2017). Ref. Green et al. (2021) observed significant discretization effects for the H dibaryon. We do not expect this to be the source of discrepancy, however. The H dibaryon is a special case, and in the $SU(3)$ flavor limit, it is expected to have a purely attractive interaction Inoue et al. (2011). Second, preliminary results for the di-nucleon on the same ensembles as the H dibaryon show milder discretization effects as well as a trend to move further from being a bound state as the continuum limit is approached Green et al. (2023).
We conclude that di-nucleons do not form bound states at heavy pion mass. In the case of the deuteron, the QC2 analysis rules out a bound state on this ensemble by $5\sigma$. In the case of the di-neutron, a bound state is excluded at the $\lesssim 3{\sigma}$ level. Previous identification of bound and deeply-bound di-nucleons Beane et al. (2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Wagman et al. (2017) most likely arose from a misidentification of a plateau in the effective energy arising from opposite sign contributions allowed in off-diagonal correlation functions, as first suggested by HAL QCD Iritani et al. (2016). This observation also invalidates all two- and higher-body LQCD matrix element calculations that utilized these off-diagonal correlation functions as those results relied upon the assumption of deeply dound di-nucleons to relate the finite and infinite volume matrix elements, which are in general non-trivial in magnitude, see for example Briceno et al. (2015), thus leaving the previous results with an unquantified systematic.
@@
DATA AVAILABILITY: The two point correlation functions as well as the ground state energies we extracted from them are available at OLCF through Constellation nn_ . Similarly, the $T_{1g}$ and $A_{1g}$ HAL QCD potential data are provided. Both data sets, as well as fit results are also available at NERSC via https://portal.nersc.gov/cfs/m2986/cosmon/nn_c103_2505.05547/. The full HAL QCD potential data can be made available to interested parties with $\sim 1$TB of space. The analysis software used to generate the results presented in this work can be found at the git repo: https://github.com/cosmon-collaboration/nn_c103_analysis.